OpenStudy (anonymous):

if i am trying to find y and it ends up being 0=0 what do i do? if i am trying to find y and it ends up being 0=0 what do i do? @Mathematics

OpenStudy (amistre64):

cry

OpenStudy (amistre64):

0=0 is always true so its all solutions

OpenStudy (amistre64):

or rather infinite solutions

OpenStudy (anonymous):

lol i already did last night.

OpenStudy (anonymous):

okay so if i am looking for x y and z and y is 0=0 then its x=5 z=-2 and y=infinite solutions?

OpenStudy (anonymous):

or the whole problem is infinite solutions?

OpenStudy (amistre64):

dunno, depends on how right you did the math from the original problem

OpenStudy (anonymous):

im pretty sure its all right.

OpenStudy (amistre64):

whats the original so that I can compare to it?

OpenStudy (anonymous):

solve by substitution (x-4y+z=6) (2x+5y-z=7) (2x-y-z=1)

OpenStudy (amistre64):

x- 4y+z = 6 2x+5y -z = 7 2x - y -z = 1 2x= 7-5y + z (7-5y + z) - y -z = 1 7-6y = 1 -6y = -6 y = 1 you sure you did it right?

OpenStudy (anonymous):

haha okay i did do it wrong. Thanks lol.

OpenStudy (amistre64):

x- 4y+z = 6 2x+5y -z = 7 2x - y -z = 1 x- 4+z = 6 2x+5 -z = 7 2x - 1 -z = 1 lol ... good luck with it ;)

OpenStudy (anonymous):

Thanks. :).

OpenStudy (anonymous):

okay so i am confused again!!! :(.

OpenStudy (amistre64):

lol ... what is it know?

OpenStudy (anonymous):

okay so i got y=1 then i got confused...:(. i feel soooooo freaking stupid!

OpenStudy (amistre64):

ok, we know y = 1, or at least it has to for this to work out; so lets go ahead and use it: x- 4+z = 6 2x+5 -z = 7 2x - 1 -z = 1 x+z = 10 2x -z = 2 2x -z = 2 the last 2 equations are the same so we can just use one of them right? x+z = 10 ; x = 10-z 2x -z = 2 2(10-z)-z=2 20 -2z -z=2 -3z = -18 z = 6

OpenStudy (amistre64):

using the first eq; x+z = 10; and z=6, x has to be 4

OpenStudy (amistre64):

dbl chk

OpenStudy (amistre64):

x- 4y+z = 6 4 1 6 ------------ 4 - 4 + 6 = 6 , good 2x+5y -z = 7 4 1 6 ------------- 8 + 5 - 6 =7; good 2x - y -z = 1 4 1 6 ------------ 8 - 1 - 6 = 1 ; good

OpenStudy (anonymous):

okay now i feel really stupid because that ended up being simple...lol thanks. Again. :).

OpenStudy (amistre64):

;)